# Pontus-Mpemba effects

### v1

@@ Abstract
we speak of a weak type-II PME. Examples are given by the middle yellow curve in Fig. 2(a) and by the lower yellow curve in Fig. 2(b), with a finite-time crossing point between the trace distances for both system copies.
Finally, the most elusive case, dubbed strong PME, may arise for the third case discussed above, see Eq. (6). It is realized if
$\mathcal{D}_{T}(\rho^{({\rm A})}(t_{\rm SI}),\rho_{\rm F})>\mathcal{D}_{T}(\rho_{\rm S},\rho_{\rm F}).$ (10)
Examples for the strong PME are given by the upper yellow curves in Figs. 2(a,b) and by the yellow curve in Fig. 2(c). This is the hardest and most counterintuitive PME to achieve, but it is closest in spirit to the Pontus example by Aristotle [39]. From an application point of view, however, we expect that the two weak PME cases will provide faster relaxation times toward F in practice.
Two-level systems.—As probably simplest application, we now turn to Markovian two-level systems. The results in Figs. 1 and 2 have been obtained for this case and show that all three PME classes can already be realized for $N=2$. For technical details and the parameter values in Figs. 1 and 2, see the End Matter section. For $N=2$, we express the traceless operators $F_{n}$ in the Lindblad equation Eq. (1) by the three Pauli matrices $\sigma_{n=1,2,3}$, and the system state is $\rho(t)=\frac{1}{2}(\mathbb{I}+\sum_{n}r_{n}(t)\sigma_{n})$ with the Bloch vector ${\bf r}(t)$. The trace distance (3) between two states $\rho_{1}$ and $\rho_{2}$ with Bloch vectors ${\bf r}_{1}$ and ${\bf r}_{2}$ is then given by ${\cal D}_{T}(\rho_{1},\rho_{2})=\frac{1}{2}|{\bf r}_{1}-{\bf r}_{2}|$. Furthermore, in Eq. (1), the system Hamiltonian is $H=\sum_{n}h_{n}\sigma_{n}$ with a real-valued vector ${\bf h}=(h_{1},h_{2},h_{3})^{T}$, while the Kossakowski matrix has the general form $C=\left(\begin{array}[]{ccc}C_{11}&C_{12}&C_{31}^{*}\\ C_{12}^{*}&C_{22}&C_{23}\\ C_{31}&C_{23}^{*}&C_{33}\end{array}\right),$ with real-valued diagonal elements and three independent complex-valued off-diagonal entries. To visualize the system dynamics within the Bloch unit ball, the velocity field $\dot{\rho}$ can be written in Bloch representation in terms of the vector field ${\bf v}({\bf r})=\dot{{\bf r}}$, where the Lindblad equation gives
@@
The $3\times 3$ matrix ${\bf\Lambda}$ and the vector ${\bf b}$ are expressed in terms of $C$ and ${\bf h}$ (see End Matter). Using these equations, we obtain Figs. 1 and 2 and the corresponding PME types. We note that for $N>2$, similar representations with ${\bf v}=\dot{\bf r}$ apply by using the $(N^{2}-1)$-dimensional coherence vector ${\bf r}$ generalizing the Bloch vector [47, 48].
As demonstrated above, all three PME classes could be realized experimentally in Markovian two-state systems. There are various physical platforms to achieve this goal with reservoir engineering techniques, e.g., trapped ions [34, 35], superconducting qubits [49], or ultracold atoms [50]. Using the specific Kossakowski matrices $C$ in the End Matter section, the jump operators $L_{j}$ in Eq. (2) follow by diagonalizing $C=U^{\dagger}\Lambda U$, where the unitary matrix $U$ contains the eigenvectors and the diagonal matrix $\Lambda$ the eigenvalues, as $L_{j}=\sum_{i}U_{ji}\sigma_{i}$. The corresponding transitions rates are given by $\gamma_{j}=\Lambda_{jj}$. The time-dependent trace distance can then be measured by quantum state tomography techniques. We conclude that an experimental realization of the PME and a test of our classification into three separate PME types seems in direct reach.
Conclusions.—We have introduced the PME as a generalized protocol realizing Mpemba physics by taking into account the time needed to prepare the initially “far” state. This protocol is more demanding but also more stringent than the standard Mpemba protocol. We find three different PME classes for Markovian systems, which have been demonstrated explicitly for two-state quantum systems. We hope that our work will motivate future experimental and theoretical work investigating and possibly extending the protocol suggested here.

### v2

@@ Abstract
we speak of a weak type-B PME. Examples are given by the middle yellow curve in Fig. 2(a) and by the lower yellow curve in Fig. 2(b), with a finite-time crossing point between both trace distances. Finally, the most elusive scenario, dubbed strong PME, may arise for the third case above, see Eq. (5), if $\mathcal{D}_{T}(\rho^{({\rm A})}(t_{\rm SI}),\rho_{\rm F})>\mathcal{D}_{T}(\rho_{\rm S},\rho_{\rm F}).$ Examples for the strong PME are given by the upper yellow curves in Figs. 2(a,b) and by the yellow curve in Fig. 2(c). This is the hardest and most counterintuitive PME to achieve, but it is closest in spirit to the Pontus example by Aristotle [40]. From an application point of view, however, we expect that the two weak PME cases will provide faster relaxation times toward F.
Two-level systems.—As probably simplest application, we now turn to Markovian two-level systems. The results in Figs. 1 and 2 were obtained for this case and show that all three PME classes can already be realized for $N=2$. For details and the parameters in Figs. 1 and 2, see End Matter. For $N=2$, we express the traceless operators $F_{n}$ in Eq. (1) by the Pauli matrices $\sigma_{n=1,2,3}$, and the state is $\rho(t)=\frac{1}{2}(\mathbb{I}+\sum_{n}r_{n}(t)\sigma_{n})$ with the Bloch vector ${\bf r}(t)$. The trace distance between $\rho_{1}$ and $\rho_{2}$ with Bloch vectors ${\bf r}_{1}$ and ${\bf r}_{2}$ is then given by ${\cal D}_{T}(\rho_{1},\rho_{2})=\frac{1}{2}|{\bf r}_{1}-{\bf r}_{2}|$. Furthermore, in Eq. (1), the Hamiltonian is $H=\sum_{n}h_{n}\sigma_{n}$ with a real-valued vector ${\bf h}=(h_{1},h_{2},h_{3})^{T}$ and the Kossakowski matrix $C=\left(\begin{array}[]{ccc}C_{11}&C_{12}&C_{31}^{*}\\ C_{12}^{*}&C_{22}&C_{23}\\ C_{31}&C_{23}^{*}&C_{33}\end{array}\right),$ with real-valued diagonal elements and three independent complex-valued off-diagonal entries. The velocity field $\dot{\rho}$ is written in Bloch representation as vector field ${\bf v}({\bf r})$, where the Lindblad equation gives
@@
The $3\times 3$ matrix $\bf\Lambda$} and the vector $\bf b$} are expressed in terms of $C$ and $\bf h$}. Using these equations, we obtain Figs. 1 and 2 and the corresponding PME types. For $N>2$, similar representations with ${\bf v}=\dot{\bf r}$ apply by using the $(N^{2}-1)$-dimensional coherence vector $\bf r$} generalizing the Bloch vector [52, 53]. All three PME classes could thus be realized experimentally in Markovian two-state systems. There are various physical platforms to achieve this goal with reservoir engineering techniques, e.g., trapped ions [35, 36], superconducting qubits [54], or ultracold atoms [55]. Using the $C$ matrices in the End Matter, by diagonalizing $C=U^{\dagger}\Lambda U$, where the unitary $U$ contains the eigenvectors and the diagonal matrix $\Lambda$ the eigenvalues, the jump operators follow as $L_{j}=\sum_{i}U_{ji}\sigma_{i}$ with rates $\gamma_{j}=\Lambda_{jj}$ in Eq. (2). The time-dependent trace distance can be measured by quantum state tomography. We conclude that an experimental realization of the PME and a test of our classification into three separate PME types seems in direct reach.
Conclusions.—We introduced the PME as a generalized Mpemba protocol taking into account the time needed to prepare the initially “far” state. It is conceptually simpler and advantageous for applications compared to the single-step Mpemba protocol. We find three PME classes which have been illustrated explicitly for Markovian two-state quantum systems. We note that non-Markovian effects [56, 29] can help to further speed up the PME relaxation dynamics, see also Ref. [57]. Appropriate distance functions could be constructed by resorting to quantum resource theories [58]. We leave this point to future work, as well as the systematic optimization of the states $\bf A$} and $\bf I$} by means of optimal control theory [44, 45, 46] and the generalization to multi-step protocols, e.g., for a study of optimal heat engines. Let us also note that for large system space dimensions, additional simplifications may be necessary to reduce the computational demands.
@@ Acknowledgements.
Data availability.—Numerical data used for generating the figures are available at the Zenodo site https://doi.org/10.5281/zenodo.17143780